OpenStudy (anonymous):

What is the value of the expression when a = 6, b = 4, and c = 8? 2a/3b-c

OpenStudy (anonymous):

is that a fraction or a division problem

OpenStudy (anonymous):

fraction

OpenStudy (anonymous):

lol i just like messin wit Freshi >:3

OpenStudy (anonymous):

i think it may be -7

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{ 2(6) }{3(4)-8 }\]

OpenStudy (anonymous):

\[\frac{ 12 }{ 12-8 }\]

OpenStudy (anonymous):

so the answer is actually 12/4

OpenStudy (anonymous):

\[\frac{ 12 }{ 4 }\]

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

3

OpenStudy (anonymous):

12/4 converts into 3

OpenStudy (anonymous):

and you would have to change it to a mixed number because it is an improper fraction

OpenStudy (anonymous):

so my answer is 12/4

OpenStudy (anonymous):

or it could be l which ever one you pick to use

OpenStudy (anonymous):

Is it asking for a fraction or?

OpenStudy (anonymous):

no it is not @corruptedalice96

OpenStudy (anonymous):

Then You should put 3, only because it seems more proper

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

why is your profile picture hunter geurink

OpenStudy (anonymous):

just asking lol

OpenStudy (anonymous):

because he is a lier

OpenStudy (anonymous):

but you have a tumblr boy just saying

OpenStudy (anonymous):

but thats fine